18 of 19 For the following electrochemical cell Sn(s)|Sn2 (aq, 0.0155 M)lIAg(aq, 3.50 M)JAg(s) Num write the net cell equation. Phases are optional. Do not include the concentrations. rour 2Ag+ + Sn → Sn+2+2Ag Calculate the following values at 25.0 ℃ using standard potentials as needed. Number Number kJ/ mol Number Number kJ/ mol Incorrect where n (sometimes symbolized as ve) is the number of electrons transferred and F is the Faraday constant, F # 96485 C/mol e. In this case, n = 2, Be sure to convert the answer from joules to kilojoules. O Previous ⓧGive Up & View Solution # Try Again Next Exit I have figured out everything except the two delta G with the red x's by them. Please help
Answer #1

The complete cell recation is as follows:

Sn + 2Ag+ -----> Sn2+ + 2Ag(s)

E°cell=E°+E°

Sn(s) ---> Sn2+ +2e- = +0.14

Ag+ + 2e- ---> Ag(s) =+0.79

E°cell= 0.93 V

now we can determine the Ecell as follows:

Ecell = E°cell - (RT/nF)ln(Q)

R=8.3145 J(/mol)(K)
T= 25C or 298K
n= 2 (2e-)
F= 96,485 c/mol ( F is a constant)
Q= products/reactants

= (0.0155)/(3.50)^2

Q= 1.26*10^-3

Now put all the values in the following equation:

Ecell = E°cell - (RT/nF)ln(Q)

Ecell= 0.93- 8.3145(298)/2(96,485)ln 1.26*10^-3

Ecell= 0.93+ 0.085
Ecell= 1.015 V

Now calculate the G°rxn as follows:

G°rxn = -nFE°cell

G°rxn = -2(96,485)(0.93)
G°rxn = -179462.1 J/mol or -179.5 KJ/ mol

Grxn = -nFEcell

Grxn = -2(96,485)(1.015)

Grxn = -195864.55J/mol or -195.86 KJ/mol